# MTU: The Multifunction Tree Unit in zkSpeed for Accelerating HyperPlonk

### v1

@@ Abstract
Zero-Knowledge Proofs (ZKPs) are critical for privacy preservation and verifiable computation. Many ZKPs rely on kernels such as the SumCheck protocol and Merkle Tree commitments, which enable their security properties. These kernels exhibit balanced binary tree computational patterns, which enable efficient hardware acceleration. Prior work has investigated accelerating these kernels as part of an overarching ZKP protocol; however, a focused study of how to best exploit the underlying tree pattern for hardware efficiency remains limited. We conduct a systematic evaluation of these tree-based workloads under different traversal strategies, analyzing performance on multi-threaded CPUs and a hardware accelerator, the Multifunction Tree Unit (MTU). We introduce a hardware-friendly Hybrid Traversal for binary tree that improves parallelism and scalability while significantly reducing memory traffic on hardware. Our results show that MTU achieves up to 1478×1478$\times$1478 × speedup over CPU at DDR-level bandwidth and that our hybrid traversal outperforms as standalone approach by up to 3×3$\times$3 ×. These findings offer practical guidance for designing efficient hardware accelerators for ZKP workloads with binary tree structures.
1 Introduction
@@
Zero-Knowledge Proofs are cryptographic protocols that allow a prover to convince a verifier that a statement is true without revealing any additional information beyond the validity of the statement itself. ZKPs have emerged as a critical building block for privacy preservation and verifiable computation in both academia and industry. They have been widely adopted in privacy-focused blockchain platforms, such as Zcash [4] and Mina [8], and are increasingly explored in applications such as privacy-preserving machine learning [18]. As real-world deployments demand better performance and scalability, hardware-accelerated implementations have become increasingly important.
@@
A key computational primitive in many modern ZKP protocols is SumCheck [36]. SumCheck plays a central role in a variety of ZKPs, including HyperPlonk [11], Spartan [33], Hyrax [37], and Libra [41]. In order to support SumCheck in the HyperPlonk protocol, several subroutines are required: Build MLE, MLE Evaluation, Multiplication Tree, and Product MLE. These workloads are detailed in Section 3.1. These subroutines share a common computational pattern: they can all be expressed as balanced binary trees, where each internal node is a computation derived from its two children. In addition to these SumCheck-related operations, Merkle Tree commitments are another binary-tree-based primitive frequently used in ZKPs, blockchains, and secure storage systems, where hash-based commitments are built in hierarchical tree structures.
@@
Balanced binary trees can be computed using either breadth-first (BFS) or depth-first (DFS) traversal strategies. While both are equivalent in algorithmic complexity, their behavior under hardware execution differs significantly due to resource constraints and memory hierarchy considerations. Prior theoretical research has primarily focused on aspects such as asymptotic complexity and software implementation [17]. However, designing hardware for these workloads requires careful consideration of several practical factors, including the ability to effectively exploit parallelism, memory traffic/bandwidth pressure, the complexity of control logic, the initiation interval in pipeline stages, and the ability to interface efficiently with upstream and downstream modules. These considerations are critical for achieving high performance but are relatively less studied in the binary tree workloads.
@@
In this paper, we systematically evaluate binary tree workloads commonly used in ZKP, especially in SumCheck-related operations. We begin by analyzing the performance scalability of different traversal methods on the multi-threaded CPU. For hardware evaluation, we utilize the Multifunction Tree Unit (MTU) architecture [13] to assess the runtime and speedup of different traversal methods, as well as to explore the design space trade-offs between area and performance across a range of hardware configurations. Additionally, we analyze the hardware-friendly Hybrid traversal scheme that balances the benefits of both BFS and DFS, enabling efficient binary tree computation on MTU-based accelerators. Our key contributions are as follows:
@@
  • We perform a systematic hardware evaluation across memory bandwidth settings ranging from DDR to HBM, and complete a design space exploration of MTU. Our results show that MTU achieves a speedup of 1478×1478$\times$1478 × over the CPU baseline with DDR-level bandwidth and that our Hybrid traversal outperforms BFS by nearly 3×3$\times$3 ×.
@@
2 Background
2.1 ZKP and zkSNARKs
ZKPs are a privacy-preserving technology that allow a prover to convince a verifier that they know some information such that a statement is true. A similar application field is private ML [27, 20, 15, 43, 16]. The state-of-the-art ZKPs are defined as Zero-Knowledge Succinct Non-interactive ARgument of Knowledge (zkSNARKs), which offer small proofs and fast verifier time at high prover costs.
zkSNARKs are constructed from the combination of an interactive oracle proof (IOP) and a polynomial commitment scheme (PCS). The IOP provides a mechanism for the prover to encode a computation as an arithmetic circuit, and subsequently, as a set of polynomials whose evaluations can be efficiently checked through queries by the verifier. SumCheck [36] is a key kernel present in many recent IOPs due to its ability to scale efficiently with increasing problem complexity. Number-Theoretic Transform (NTT) [44] is an alternative to SumCheck used in some ZKP protocols. The PCS, in turn, binds the prover to these polynomials and enables the verifier to check the correctness without requiring access to the full polynomials. There are two common PCS kernels, Multi-Scalar Multiplication (MSM) and Merkle Tree. Together, IOPs and PCSs form the foundation of zkSNARK protocols such as HyperPlonk [11] and Orion[42]. We focus on workloads related to SumCheck and Merkle Tree, seen in HyperPlonk and Orion respectively, as they both contain binary tree computational patterns and can be accelerated using similar hardware.
2.2 SumCheck and MLEs
SumCheck [36] is an interactive protocol that allows a prover to convince a verifier that the sum of a multivariate polynomial, $f(x_{1},\ldots,x_{\mu})$, over the boolean hypercube is some claimed sum, $S$:
$S=\sum_{x_{1}\in\{0,1\}}\sum_{x_{2}\in\{0,1\}}\cdots\sum_{x_{\mu}\in\{0,1\}}f(x_{1},x_{2},\dots,x_{\mu})$ (1)
The boolean hypercube is equivalent to a $\mu$-bit space $\vec{x}\in\{0,1\}^{\mu}$. At a high level, for a $\mu$-variable polynomial $f$, there are $\mu$ rounds of the protocol; in round $i$, the prover makes a claim about the variable $x_{i}$. In zkSNARK protocols like HyperPlonk, a key constraint about $f$ is that it is multilinear, meaning each variable’s maximum degree in $f$ is one. A general representation of multilinear polynomials, for example, with $\mu=3$,
$f(x_{1},x_{2},x_{3})=f(0,0,0)(1-x_{1})(1-x_{2})(1-x_{3})\leavevmode\nobreak\ +$ (2)
@@ 2.2 SumCheck and MLEs
Throughout SumCheck rounds, $f$ will be evaluated at points where the variables $x_{i}$ take values that are not limited to 00 or $1$. Even though $f$ is originally defined (and stored in hardware) over Boolean hypercube, the values of $f$ at non-Boolean inputs will be extrapolated based on the values stored in the lookup table. These resulting values are called extensions, thus $f$ is called a multilinear extension (MLE). The term MLE table denotes the corresponding lookup table structure mentioned above, where each entry can occupy 256 bits in ZKP applications.
@@
2.3 Merkle Tree Commitment
Figure 1: Dataflow for computing the Build MLE workload (output size $2^{5}$) using a binary tree traversal. The result level (blue) is indexed as Level 1 and consists of $2^{5}$ entries. An example partitioning into four subtrees is highlighted with red dash-boxes to illustrate parallel DFS processing.
@@ 2.3 Merkle Tree Commitment
Merkle tree in ZKP allow a prover to commit to a large vector of data using a compact root hash. Each leaf node of the tree corresponds to a hashed element in the original vector, while each internal node is the cryptographic hash (e.g., SHA3 [6], Poseidon Hash [21]) of its two child nodes. This recursive construction culminates in a single root hash that serves as a commitment to the entire vector. Unlike MSM-based PCS constructions, Merkle trees do not require elliptic curve operations and thus typically offer faster prover runtime. However, this comes at the cost of increased proof sizes and slower verifier performance. Merkle-based PCS protocols, such as Orion [42], leverage this tradeoff by targeting environments where prover speed is critical and verifier costs are amortized or less constrained. Constructing a Merkle tree naturally maps to a binary tree reduction pattern. Similar to SumCheck, Merkle tree construction benefits from an efficient tree unit, particularly when building large trees over vectors with lengths $\geq 2^{20}$.
3 Binary Tree Traversal
3.1 Binary Tree Workloads in ZKP
@@
Balanced binary tree computations are prevalent in ZKP protocols, particularly in SumCheck-based constructions [11, 33]. These computation patterns appear repeatedly across ZKP applications and often form performance bottlenecks. Fortunately, their shared structure allows the design of a general hardware module that can accelerate them all with minimal specialization. zkSpeed introduces the MTU as a module capable to support multiple binary-tree-based computations. In this section, we highlight several common binary tree computations in ZKP that motivate the design of the MTU.
3.1.1 Build MLE
@@
Build MLE is an important subroutine in accelerating SumCheck-related protocols. SumCheck allows the prover to prove the sum of a multivariate polynomial $f$ over the Boolean HyperCube is zero, but in many ZKP protocols, the goal could be stronger: to prove that $f(\vec{x})=0$ for every Boolean HyperCube input, i.e., each lookup table entry evaluates to 0. However, running a SumCheck directly on $f(\vec{x})$ with a claimed sum $S=0$ is not sufficient, because two terms might be nonzero yet cancel each other out, meaning the total sum is zero even if some of the individual evaluations are not zero.
To ensure correctness, a polynomial $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ is introduced [7, 11, 33, 34]:
$\tilde{\mathrm{eq}}(\vec{x},\vec{r})=\prod_{i=1}^{\mu}(r_{i}x_{i}+(1-r_{i})(1-x_{i}))$ (3)
@@
where $\vec{x}\in\{0,1\}^{\mu},\vec{r}\in{{\mathbb{F}}}^{\mu}$ are $\mu$ random values in ${{\mathbb{F}}}$ finite field (challenges). By proving $\sum_{\vec{x}\in\{0,1\}^{\mu}}(f(\vec{x})\cdot\tilde{\mathrm{eq}}(\vec{x},\vec{r}))=0$ with SumCheck, it implies that $f(\vec{x})=0$ $\forall\vec{x}\in\{0,1\}^{\mu}$. Its correctness is shown in [33, 10] and is known as a ZeroCheck in [19, 11]. We call the computation to construct $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ as Build MLE.
@@
To perform ZeroCheck, the prover must evaluate $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ at all $\vec{x}\in\{0,1\}^{\mu}$, which in fact are a list of products from $(1-r_{1})(1-r_{2})\cdots(1-r_{\mu})$ to $r_{1}r_{2}\cdots r_{\mu}$. A naive approach requires $(\mu-1)2^{\mu}$ modular multiplications to compute all such products individually. However, this cost can be reduced substantially by computing the values via a binary tree structure, as illustrated in Figure 1. The binary tree method reduces the number of modular multiplications to $42^{\mu+1}-$ 4, providing significant computational savings. This optimization motivates the tree-based hardware design capable of efficiently supporting this pattern. In the tree-based computation, we can reduce one modular multiplication by using
$(1-r_{1})\cdot(1-r_{2})=(1-r_{1})-(1-r_{1})\cdot r_{2}$ (4)
That is, $(1-r_{1})(1-r_{2})$ can be computed from $(1-r_{1})r_{2}$, instead of computing an extra modular multiplication with $(1-r_{2})$ and $(1-r_{1})$.
@@
Figure 2: Dataflow for computing the MLE evaluation workload of input size $2^{4}$. The input level is indexed as Level 1, and the final output is highlighted in blue. An example partitioning into two subtrees is shown with red dash-boxes to illustrate parallel DFS processing.
3.1.2 MLE Evaluation
@@
$f(r_{1},\cdots,r_{\mu})=f(0,\cdots,0)(1-r_{1})\cdots(1-r_{\mu})+\cdots$ (5)
$+\leavevmode\nobreak\ f(1,\cdots,1)x_{1}\cdots x_{\mu}$
A naive method is to modular-multiply every individual term and sum all at the end. Similar to Build MLE, we can substantially improve the computation by following the binary tree structure, while the tree is now inverted (reduction computation, as shown in Figure 2). With the binary tree structure, the number of modular multiplications can be optimized from $\mu\cdot 2^{\mu}$ to $12^{\mu}-$ 1.
Note that in Figure 2 we also applied this trick to further reduce one modular multiplication.
$f_{0}\cdot(1-r_{1})+f_{1}\cdot r_{1}=f_{0}+r_{1}\cdot(f_{1}-f_{0})$ (6)
Since modular addition is cheaper than modular multiplication, we can replace the original form on the left of the equation with only one mod-mul on the right.
3.1.3 Merkle Tree
The Merkle tree is a widely used cryptographic commitment scheme in which the root of a binary tree serves as the commitment to a set of leaf values. Individual leaves can later be revealed and efficiently verified as part of the committed structure using authentication paths. In the context of ZKP protocols, Merkle trees are often employed as hash-based polynomial commitment schemes.

### v2

# MTU: The Multifunction Tree Unit for Accelerating Zero-Knowledge Proofs

@@ Abstract.
Zero-Knowledge Proofs (ZKPs) are critical for privacy-preserving techniques and verifiable computation. Many ZKP protocols rely on key kernels such as the SumCheck protocol and Merkle Tree commitments to enable their key security properties. These kernels exhibit balanced binary tree computational patterns, which enable efficient hardware acceleration. Although prior work has investigated accelerating these kernels as part of an overarching ZKP protocol, exploiting this common tree pattern remains relatively underexplored. We conduct a systematic evaluation of these tree-based workloads under different traversal strategies, analyzing performance on multi-threaded CPUs and the Multifunction Tree Unit (MTU) hardware accelerator. We introduce a hardware-friendly Hybrid Traversal for binary tree that improves parallelism and scalability while significantly reducing memory traffic on hardware. Our results show that MTU achieves up to $1478\times$ speedup over CPU at DDR-level bandwidth and that our hybrid traversal outperforms breadth-first search by up to $3\times$. These findings offer practical guidance for designing efficient hardware accelerators for ZKP workloads with binary tree structures.
zero-knowledge proofs, hardware acceleration, tree traversal
journalyear: 2025copyright: rightsretainedconference: Hardware and Architectural Support for Security and Privacy 2025; October 19, 2025; Seoul, Republic of Koreabooktitle: Hardware and Architectural Support for Security and Privacy 2025 (HASP 2025), October 19, 2025, Seoul, Republic of Koreadoi: 10.1145/3768725.3768737isbn: 979-8-4007-2198-4/25/10ccs: Computer systems organization Special purpose systemsccs: Security and privacy Cryptography
1. Introduction
@@
Zero-Knowledge Proofs are cryptographic protocols that allow a prover to convince a verifier that a statement is true without revealing any additional information beyond the validity of the statement itself. ZKPs have emerged as a critical building block for privacy-preserving techniques and verifiable computation in both academia and industry. They have been widely adopted in privacy-focused blockchain platforms, such as Zcash (Ben-Sasson et al., 2014) and Mina (Bonneau et al., 2020), and are increasingly explored in applications such as privacy-preserving machine learning (Feng et al., 2023; Mo et al., 2023a). As real-world deployments demand better performance and scalability, hardware-accelerated implementations have become increasingly important.
@@
A key computational primitive in many modern ZKP protocols is the SumCheck protocol (Thaler, 2022). SumCheck plays a central role in a variety of ZKPs, including HyperPlonk (Chen et al., 2022), Spartan (Setty, 2020), Hyrax (Wahby et al., 2017), and Libra (Xie et al., 2019). In order to support SumCheck in the HyperPlonk protocol, several subroutines are required: Build MLE, MLE Evaluation, Multiplication Tree, and Product MLE. These workloads are detailed in Section 3.1. These subroutines share a common computational pattern: they can all be mapped onto balanced binary trees, where each internal node is a computation derived from its two children. In addition to these SumCheck-related operations, Merkle Tree commitments are another binary-tree-based primitive frequently used in ZKPs and other cryptographic systems, where hash-based commitments are built in hierarchical tree structures.
@@
Balanced binary trees can be computed using either breadth-first (BFS or level-ordered) or depth-first (DFS) strategies. While both are equivalent in algorithmic complexity, their behavior under hardware execution differs significantly due to resource constraints and memory hierarchy considerations. Prior research has primarily focused on theoretical aspects such as asymptotic complexity and software implementation (Erickson, 2023). However, designing hardware for these workloads requires careful consideration of several practical factors, including parallelism efficiency, memory traffic/bandwidth pressure, the complexity of control logic, the initiation interval in pipeline stages, and the ability to interface efficiently with upstream and downstream modules. These considerations are critical but relatively less studied in the binary tree workloads.
@@
In this paper, we address this gap by systematically evaluating binary tree workloads under various traversal strategies. We analyze the performance scalability of different traversal methods on CPU. For hardware evaluation, we utilize the Multifunction Tree Unit (MTU) architecture (Daftardar et al., 2025b, a) to assess the runtime and speedup of different traversal methods, and explore the design space trade-offs between area and performance across a range of hardware configurations. The MTU is integrated in zkSpeed (Daftardar et al., 2025b) to perform end-to-end proof generation. Additionally, we analyze the hardware-friendly Hybrid traversal scheme that balances the benefits of both BFS and DFS, enabling efficient binary tree computation on MTU-based accelerators. Our contributions are as follows:
@@
  • We perform a systematic hardware evaluation across memory bandwidth settings ranging from DDR to HBM, and complete a design space exploration of MTU. Our results show that MTU achieves a speedup of $1478\times$ over the CPU baseline with DDR-level bandwidth and that our Hybrid traversal outperforms BFS by nearly $3\times$.
@@
2. Background
2.1. ZKP and zkSNARKs
ZKPs are a privacy-preserving technology that allow a prover to convince a verifier that they know some information such that a statement is true. The state-of-the-art ZKPs are Zero-Knowledge Succinct Non-interactive ARgument of Knowledge (zkSNARK), which offer small proofs and fast verifier time at high prover costs.
zkSNARKs are constructed from the combination of an interactive oracle proof (IOP) and a polynomial commitment scheme (PCS). The IOP provides a mechanism for the prover to encode a computation as an arithmetic circuit, and subsequently, as a set of polynomials whose evaluations can be efficiently checked through queries by the verifier. SumCheck is a key kernel present in many recent IOPs due to its ability to scale efficiently with increasing problem complexity. Number-Theoretic Transform (NTT) is an alternative to SumCheck used in some ZKP protocols. The PCS, in turn, binds the prover to these polynomials and enables the verifier to check the correctness without requiring access to the full polynomials. There are two common PCS kernels, Multi-Scalar Multiplication (MSM) and Merkle Tree. Together, IOPs and PCSs form the foundation of zkSNARK protocols such as HyperPlonk (Chen et al., 2022) and Orion(Xie et al., 2022). We target workloads related to SumCheck and Merkle Tree, seen in HyperPlonk and Orion respectively, due to their shared binary tree computation patterns suited for similar hardware acceleration.
2.2. SumCheck and MLEs
SumCheck (Thaler, 2022) is an interactive protocol that allows a prover to convince a verifier that the sum of a multivariate polynomial, $f(x_{1},\ldots,x_{\mu})$, over the boolean hypercube is some claimed sum, $S$: $S=\sum_{x_{1}\in\{0,1\}}\sum_{x_{2}\in\{0,1\}}\cdots\sum_{x_{\mu}\in\{0,1\}}f(x_{1},x_{2},\dots,x_{\mu})$. The boolean hypercube is equivalent to a $\mu$-bit space $\vec{x}\in\{0,1\}^{\mu}$. At a high level, for a $\mu$-variable polynomial $f$, there are $\mu$ rounds of the protocol; in round $i$, the prover makes a claim about the variable $x_{i}$. In protocols like HyperPlonk, a key constraint about $f$ is that it is multilinear, meaning each variable’s maximum degree in $f$ is one. A general representation of multilinear polynomials, for example, with $\mu=3$,
(1) $f(x_{1},x_{2},x_{3})=f(0,0,0)(1-x_{1})(1-x_{2})(1-x_{3})~+$
@@ 2.2. SumCheck and MLEs
Throughout SumCheck rounds, $f$ will be evaluated at non-Boolean values of $x_{i}$, even though $f$ is originally defined (and stored in hardware) over Boolean inputs. This is done by extrapolating from values stored in the lookup table. These values are called extensions, thus $f$ is called a multilinear extension (MLE). The term MLE table denotes the corresponding lookup table structure mentioned above, where each entry’s data value (typically) occupies  256 bits in ZKP applications.
@@
2.3. Merkle Tree Commitment
@@
Merkle trees in ZKP allow a prover to commit to a large vector of data using a compact root hash. Each leaf node of the tree corresponds to a hashed element in the original vector, while each internal node is the cryptographic hash (e.g., SHA3 (Bertoni et al., 2009), Poseidon Hash (Grassi et al., 2019)) of its two child nodes. This recursive construction culminates in a single root hash that serves as a commitment to the entire vector. Unlike MSM-based PCS constructions, Merkle trees do not require elliptic curve operations and thus typically offer faster prover runtime. However, this comes at the cost of increased proof sizes and longer verifier latency. Merkle-based PCS protocols, such as Orion (Xie et al., 2022), leverage this tradeoff by targeting environments where prover speed is critical and verifier costs are amortized or less constrained. Merkle tree construction naturally maps to a binary tree reduction pattern, which benefits from an efficient tree hardware unit, especially for vectors of lengths $\geq 2^{20}$.
3. Binary Tree Traversal
3.1. Binary Tree Workloads in ZKP
@@
Balanced binary tree computations are prevalent in ZKP protocols, particularly in SumCheck-based constructions (Chen et al., 2022; Setty, 2020). These computation patterns appear repeatedly across ZKP applications and often form performance bottlenecks. Fortunately, their shared structure allows the design of a general hardware module that can accelerate them all with minimal specialization. zkSpeed introduces the MTU as a module capable to support multiple binary-tree-based computations. In this section, we highlight several common binary tree computations in ZKPs that motivate the design of the MTU.
3.1.1. Build MLE
@@
Build MLE is an important subroutine in SumCheck related protocols. SumCheck allows the prover to prove the sum of a multivariate polynomial $f$ over the Boolean HyperCube is zero, but in many ZKP protocols, the goal could be stronger: to prove that $f(\vec{x})=0$ for every Boolean HyperCube input, i.e., each lookup table entry evaluates to 0. However, running a SumCheck directly on $f(\vec{x})$ with a claimed sum $S=0$ is not sufficient, because two terms might be nonzero yet cancel each other out, meaning the total sum is zero even if individual evaluations are not zero.
@@
To ensure correctness, a polynomial $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ is introduced (Blumberg et al., 2014; Chen et al., 2022; Setty, 2020; Setty and Lee, 2020):   $\tilde{\mathrm{eq}}(\vec{x},\vec{r})=\prod_{i=1}^{\mu}\big(r_{i}x_{i}+(1-r_{i})(1-x_{i})\big)$,   where $\vec{x}\in\{0,1\}^{\mu}$, and $\vec{r}\in{{\mathbb{F}}}^{\mu}$ is a $\mu$-length vector of random values (challenges) in a finite field $\mathbb{F}$}}. By proving $\sum_{\vec{x}\in\{0,1\}^{\mu}}\big(f(\vec{x})\cdot\tilde{\mathrm{eq}}(\vec{x},\vec{r})\big)=0$ with SumCheck, it implies that $f(\vec{x})=0$ $\forall\vec{x}\in\{0,1\}^{\mu}$. Its correctness is shown in (Setty, 2020; Campanelli et al., 2019) and is known as a ZeroCheck in (Gabizon et al., 2019; Chen et al., 2022). We denote the computation of $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ as Build MLE.
@@
To perform ZeroCheck, the prover must evaluate $\tilde{\mathrm{eq}}(\vec{x},\vec{r})$ at all $\vec{x}\in\{0,1\}^{\mu}$, which in fact is a list of products from $(1-r_{1})(1-r_{2})\cdots(1-r_{\mu})$ to $r_{1}r_{2}\cdots r_{\mu}$. A naive approach requires $(\mu-1)2^{\mu}$ modular multiplications to compute all such products individually. However, this cost can be reduced substantially by computing the values via a binary tree structure, as illustrated in Figure 1. The binary tree method reduces the number of modular multiplications to $2^{\mu+1}-4$, providing significant computational savings. This optimization motivates the tree-based hardware design capable of efficiently supporting this pattern. In the tree-based computation, we can reduce one modular multiplication by using: $(1-r_{1})\cdot(1-r_{2})=(1-r_{1})-(1-r_{1})\cdot r_{2}$. That is, $(1-r_{1})(1-r_{2})$ can be computed from $(1-r_{1})r_{2}$, instead of computing an extra modular multiplication with $(1-r_{2})$ and $(1-r_{1})$.
@@
Figure 1. Dataflow for computing Build MLE (output size $2^{5}$). The result level (blue) is indexed as Level 1. Partitioning into four subtrees is highlighted with red dashed boxes to illustrate parallel DFS processing.
3.1.2. MLE Evaluation
@@
(2) $f(r_{1},\cdots,r_{\mu})=\big(f(0,\cdots,0)(1-r_{1})\cdots(1-r_{\mu})\big)+\cdots$
$+~\big(f(1,\cdots,1)x_{1}\cdots x_{\mu}\big)$
A naive method is to modular-multiply every individual term and sum all at the end. Similar to Build MLE, we can substantially improve the computation by following the binary tree structure, while the tree is now inverted (reduction computation, as shown in Figure 2). With the binary tree structure, the number of modular multiplications can be optimized from $\mu\cdot 2^{\mu}$ to $2^{\mu}-1$. In Figure 2 we applied this trick to further reduce one modular multiplication: $f_{0}\cdot(1-r_{1})+f_{1}\cdot r_{1}=f_{0}+r_{1}\cdot(f_{1}-f_{0})$. Since modular addition is cheaper than modular multiplication, we can replace the original form on the left of the equation with one mod-mul on the right.
3.1.3. Merkle Tree
The Merkle tree is a widely used cryptographic commitment scheme in which the root of a binary tree serves as the commitment to a set of leaf values. Individual leaves can later be revealed and efficiently verified as part of the committed structure using authentication paths. In ZKP protocols, Merkle trees are often employed as hash-based polynomial commitment schemes.